OpenStudy (anonymous):

The ratio of the areas of two similar polygons is 121:225. If the perimeter of the first polygon is 60 cm, what is the perimeter of the second polygon? Round to the nearest tenth

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Im guessing 121/60: 225/x then 121x=13500 13500/121 giving me 111.570 for the first part?????

OpenStudy (anonymous):

Am I way off?

OpenStudy (skullpatrol):

Your answer is correct.

OpenStudy (anonymous):

So my answer is 111.6

OpenStudy (skullpatrol):

units please

OpenStudy (anonymous):

111.6cm :)

OpenStudy (skullpatrol):

Good work

OpenStudy (anonymous):

Thank you !

Directrix (directrix):

If two polygons are similar, the scale factor is the same as the ratio of the perimeters. A theorem states that if two polygons are similar, the square of the scale factor is the ratio of the areas of the two similar polygons.

OpenStudy (anonymous):

????

Directrix (directrix):

121/225 = (60/p)^2 where p is the perimeter of the larger polygon.

Directrix (directrix):

11/15 = 60/p Solving for p does not render 11.6. Perhaps, I made an error. @skullpatrol

OpenStudy (anonymous):

111.6 thats rounded the original is 111.570

OpenStudy (skullpatrol):

@Directrix I am unaware of this "theorem"

Directrix (directrix):

What theorem supports this statement --> 121/60: 225/x

Directrix (directrix):

Theorem: If two similar triangles have a scale factor of a : b, then the ratio of their areas is a^2 : b^2

OpenStudy (anonymous):

@Directrix I FINALLY UNDERSTOOD YOU! Thanks for dealing with me, I've marked you best answer for all of them, and I became a follower. @adillie Sorry for budding in, I just wanted to thank Directrix before I left.

Directrix (directrix):

Theorem: If two similar triangles have a scale factor of a : b, then the ratio of their perimeters is a : b.

Directrix (directrix):

The ratio of areas of two similar figures is the square of the scale factor. The ratio of volumes of two similar figures is the cube of the scale factor. http://www.northstarmath.com/sitemap/SimilarFigures.html

OpenStudy (anonymous):

Have to close this question

OpenStudy (anonymous):

81.81 was right

Directrix (directrix):

Thanks for letting me know. I did not want you to get an incorrect answer.